OpenStudy (anonymous):

Write the following as a common fraction or mixed number. Write your answer in lowest terms. 0.03

OpenStudy (anonymous):

3/100

OpenStudy (anonymous):

how did you get that answer?

OpenStudy (anonymous):

\[0.03=0.03/1*100=3/100\]